# Learning to Infer Parameterized Representations of Plants from 3D Scans

### v1

@@ Abstract
Reconstructing faithfully the 3D architecture of plants from unstructured observations is a challenging task. Plants frequently contain numerous organs, organized in branching systems in more or less complex spatial networks, leading to specific computational issues due to self-occlusion or spatial proximity between organs. Existing works either consider inverse modeling where the aim is to recover the procedural rules that allow to simulate virtual plants, or focus on specific tasks such as segmentation or skeletonization. We propose a unified approach that, given a 3D scan of a plant, allows to infer a parameterized representation of the plant. This representation describes the plant’s branching structure, contains parametric information for each plant organ, and can therefore be used directly in a variety of tasks. In this data-driven approach, we train a recursive neural network with virtual plants generated using an L-systems-based procedural model. After training, the network allows to infer a parametric tree-like representation based on an input 3D point cloud. Our method is applicable to any plant that can be represented as binary axial tree. We evaluate our approach on Chenopodium Album plants, using experiments on synthetic plants to show that our unified framework allows for different tasks including reconstruction, segmentation and skeletonization, while achieving results on-par with state-of-the-art for each task.
@@
11footnotetext: Inria centre at the University Grenoble Alpes22footnotetext: Inria centre of Lyon
@@ 1 Introduction
Inferring high-level information of plants from observations in a way that respects the biological traits (e.g. the radius of a lateral branch is never larger than that of it’s parent stem) is an important problem that has been studied for decades [52]. High-level information includes the reconstruction of the 3D geometry of a plant, the segmentation of a plant into parts, and the extraction of a skeleton or measurements. Observations are typically 2D images or 3D scans. Automatically extracting this information allows to use reconstructed 3D plants in content generation and virtual reality [29], and to use part segmentation, skeleton information, and measurements in plant phenotyping, e.g. [33, 13, 18].
@@
The problem of inferring high-level information is challenging because plants have complex structures due to their branching systems, which leads to strong self-occlusions and ambiguities. Existing works that analyze 2D or 3D plant observations therefore either focus on inverse modeling, where the goal is to find growth rules that allow to generate a given plant, or are task-specific. Inverse modeling is challenging and existing methods are limited to simple branching structures [24, 48]. Task-specific methods focus on 3D reconstruction e.g. [29], organ segmentation e.g. [33], skeletonization e.g. [13], or measurement extraction e.g. [18].
We propose a unified approach that, given a 3D scan of a plant, allows to infer a parameterized representation of the plant. This representation contains parametric information on each plant organ, and can therefore be directly used to reconstruct the 3D geometry of the plant with corresponding segmentation labels and measurements, thereby allowing to solve a variety of tasks. We achieve this by learning a shape space of 3D plants based on representations generated by a biologically inspired procedural model, in a way that for any point in shape space, we can retrieve its corresponding parameterized representation. By learning a function that maps an observed 3D point cloud to the learned shape space, our work allows to find the representation of a plant without reconstructing the full procedural model rules. Hence, our work can be considered a step towards inverse modeling-based reconstruction.
The key idea that allows our method to address the challenging problem of inferring high-level information from 3D scan data plagued by occlusions is to learn a shape space with the help of instances generated using a procedural model made of recursive rules that describe the development of the plant architecture. We leverage a Lindenmayer-systems (L-systems) model [40] allowing to generate plants, and the recursive nature of plant growth in a recursive neural network to learn a latent distribution of the plants’ geometry and branching structure. We generate a synthetic dataset using procedural modeling rules, which allows for supervised learning from large amounts of occlusion-free training data. The synthetic plants are represented in a binary axial tree form that can be geometrically interpreted to obtain the 3D virtual plant. The binary tree format is exploited using a recursive auto-encoder. To address the problem of generalizing to 3D scans plagued by acquisition noise and occlusions, we virtually simulate the acquisition noise on the 3D virtual plants, in addition to simulated monocular depth maps during training and learn a mapping from simulated 3D scan data to the shape space.
We evaluate our method on synthetic Chenopodium Album plants, an annual plant, in early growth stages. We choose this plant because of the availability of real scans in dataset [34], and its simple structure in early stage that can be encoded into a binary tree representation. We use our method to perform the different tasks of 3D reconstruction, skeleton extraction and plant part segmentation. Our experiments show results that are on-par with the state-of-the-art, with more robustness to noisy and partial inputs. We also show qualitative results for instance segmentation, and discuss the potentials to apply our method on real scans.
@@
  • A unified approach to infer a parameterized representation of a plant from a 3D scan based on a recursive neural network.
  • 3D plant reconstruction, skeletonization, organ segmentation, with performance that is on-par with the state-of-the-art on each task.
  • A dataset of virtual plants, in a form of L-Strings, that represents instances of Chenopodium Album plants in early growth stages.
@@ 2 Related Work
Vision based systems and platforms have been developed for plant phenotyping and crops analysis [7, 8, 11, 10, 37]. Furthermore, plant scan datasets and data generation methods were provided for training data-driven models, e.g.  the rosebush annotated dataset ROSE-X [15], an annotated dataset of 3 plant species PLANEST-3D [31], and a labeled point cloud generation method of virtual plants [4]. Furthermore, small datasets of 3D scans were captured e.g. [34].
Existing works on plant phenotyping that deal with plant scan data belong to two main categories: works that aim for inverse plant modeling, and task-specific methods that aim to extract information. Our work is in between these categories, since it is centered around the inference of parameterized plant representations, which allows to extract various high-level information including the full branching structure and 3D geometry of the plant. This can be seen as a step towards inverse plant modeling as we infer an instance of a procedural model.
@@ 2.1 Inverse Plant Modeling
Procedural plant models generate a plant following growth rules. On the contrary, inverse plant modeling aims to find the rules that allow to generate the given plant. In other words, the goal is to infer the model given the plant geometry as input. Guo et al. [23] and Štáva et al. [47] infer procedural modeling rules generated using an L-system from 2D images of simple branching structures. Štáva et al. [48] introduce a parametric procedural model and estimate the parameters that allow this model to generate trees of a single species similar to an input 3D mesh. More recently, Lee et al. [24] used transformers to learn a latent space that allows to predict L-system rules that correspond to 3D point clouds of a tree, and to generate similar trees.
As Lee et al. [24], we use data-driven neural networks to learn a latent representation of the plant. However, we aim to directly predict the L-String that represents the input plant with its branching structure, geometrical and other biological traits, instead of predicting the L-system rules.
@@ 2.2 Task Specific Methods
Existing methods used in plant phenotyping from 3D input point clouds mainly focus on four tasks: 3D reconstruction of the plant, extracting a skeleton of the branching structure of the plant, segmenting the plant into organs, and extracting measurements from the point cloud.
@@
Reconstructing plant geometry from an input scan represented by a 3D point cloud is an important problem due to its applications in computer graphics and plant phenotyping. This is a particularly challenging 3D reconstruction problem as plants contain thin structures and as different organs lead to significant self-occlusions.
In computer graphics, Gonzalez et al. [22] use UnrealEngine [41] to reconstruct digital urban trees by reconstructing a mesh that represents the tree trunk, estimating the volume and density of the canopy, and filling the canopy with generated leaves. This method focuses on global geometry and ignores the branching structure of the plant architecture.
Another line of work focuses on the tree branching structure [14, 39, 54, 29, 53], by estimating a skeletal structure of the tree. Given the skeleton, a mesh surface can be reconstructed over each branch. However, the foliage part of the trees is either excluded from the input, or reconstructed by randomly generating leaves to fill the canopy.
Closest to our work, Prasad et al. [38] explore different implicit surface reconstruction algorithms, including classical and neural based ones, on 3D point clouds of a single plant. They demonstrate that this task is challenging due to fine geometric structures. Note that our method’s output contains more information than a 3D reconstruction, as our parametric output allows to infer information on the plant’s skeleton and organs. For 3D reconstruction, we compare our method experimentally to this work.
@@
Extracting the skeletal structure of plants is a well-studied problem. One line of works, discussed above, extracts skeletons as one part of a pipeline to reconstruct the plant [14, 39, 54, 29, 53].
Another line of works corresponds to optimization-based methods. Meyer et al. [32] use a point contraction algorithm to extract skeletons of leafless cherry trees. Works using graph-based approaches have shown to extract skeletal structures even in the presence of noise for trees [53, 30]. A recent deep learning-based approach uses a sparse convolutional neural network to estimate the medial axis of a tree [13]. Given a point cloud of a plant with an initial extracted skeleton, Chaudhury et al. [6] propose a stochastic optimization to refine the initial skeleton by estimating point transformations.
All discussed methods require a leafless input point cloud to estimate the skeleton. When leaves are present in the input point cloud, these methods estimate skeletal structures inside of the leaves, thereby adding noise to the output. Since our method outputs a segmented parameterized representation of the plant, extracting the skeleton of a leafy input plant does not pose additional challenges. We experimentally compare our method to one successful approach from each category for which code is available, namely [53, 30, 6].
@@
Segmenting different organs from point clouds of plants is critical for phenotyping on both semantic and instance levels. Both learning and optimization-based techniques were proposed in the literature to segment a plant into organs. Mirande et al. [33] propose a graph-based optimization approach with botanical knowledge refinement for semantic and instance segmentation. Wahabzada et al. [51] present an unsupervised data-driven method using histogram clustering in a Euclidean space. For supervised learning methods, a benchmark for plant organ segmentation [49] has been released based on the ROSE-X dataset [15]. More complex deep learning architectures designed for organ segmentation lead to accurate results [26, 50, 25]. We compare our method experimentally to two of these successful approaches [26, 25].
Measurement Extraction
Extracting phenotypic measures like leaf area index or main stem height provides insight about plant and crop behavior in different conditions. Paproki et al. [36] propose a mesh processing technique for 3D plants phenotypic traits analysis from multi-view images. Chattopadhyay et al. [3] model and reconstruct apple trees with diameter estimation for the application of automatic dormant pruning. Ghahremani et al. [18] apply the RANSAC algorithm to model plant organs that fit an input point cloud to estimate phenotypic features like angles and diameters of leaves and branches.
@@
Our experiments focus on the three tasks of 3D reconstruction, skeletonization and segmentation. As shown in Table 1, only few works perform the three tasks of 3D reconstruction, skeletonization and segmentation, like [29, 54]. However, these methods are designed to be applied on big trees with trunks and foliage, and are not applicable on small annual plants where the leaf size to stem diameter proportion is different. On annual plants, our method remains generic due to the learned parametric model that contains the information needed for different tasks.
Work Recon- Skeletoni- Segment- Annual
struction zation ation plants
Gonzalez
 et al. [22]
Liu
 et al. [29]
Du
 et al. [14]
Linvy
 et al. [30]
Preuksakarn
 et al. [39]
Parsad
 et al. [38]
Dobbs
 et al. [13]
Chaudhury
 et al. [6]
Yan
 et al. [54]
Meyers
 et al. [32]
Mirande
 et al. [33]
Turgut
 et al. [49]
Turgut
 et al. [50]
Wahabzada
 et al. [51]
Li
 et al. [26]
Li
 et al. [25]
Paproki
 et al. [36]
Ghahremani
 et al. [18]
Ours
Table 1: Positioning of our method w.r.t. the ability to perform 3D reconstruction, skeletonization and part segmentation, and the applicability on small annual plants.
@@ 3 Background
We propose a method that learns a shape space of a plant species with the help of recursive neural networks. To learn this shape space, we rely on a large dataset of 3D plants with corresponding parametric representations that has been simulated with the help of plant models. This section provides background on plant modeling and on recursive neural networks.
@@ 3.1 Plant Modeling
Plant modeling aims to find a mathematical model that describes the complex geometry and growth rules for a species of plants. Moreover, plant models respect biological rules to allow realistic simulation of the plant’s appearance and behavior in different environmental conditions [5, 19]. The importance of plant modeling comes from its applications in biological and agronomical studies. The synthetic models can be used to perform growth simulation experiments, or estimate biological traits and growth patterns of plants. This allows to study plants with fast and non-invasive methods.
Procedural methods are among the most commonly used plant modeling approaches, including L-systems [40], the space colonization algorithm [43], and commercial tools such as SpeedTree. In our work, we use procedural plant models to generate data for training and evaluating our neural network.
Our implementation uses an L-system-based procedural approach [40] to generate plant data, a classical rule based plant modeling technique, that is particularly convenient due to the availability of tools such as L-Py [2]. In L-systems, a plant is represented as a string of symbols, called L-String, possibly bearing geometrical or biological parameters. Similarly to natural or computer language grammars, an L-system consists of a set of rewriting rules that define how plant components represented by L-String symbols change as time proceeds, by specifying how symbols get replaced by combinations of other symbols [40, 21]. At each step, the rules replace the symbols in parallel, resulting in a new L-String representing the next plant state.
L-Strings of plants mainly consist of two types of symbols: modules and brackets. Modules can refer to various plant parts e.g. stems, leaves, flowers, etc. Opening and closing brackets indicate the start and the end of every branch in the plant. Two successive modules in an L-sting have a parent-child relationship. Open brackets allow for a module to have more than one child, and for different children to have siblings relationships. However, each parent has at most one special child that corresponds to its successor on the same plant axis. Such L-Strings encode axial trees [40] that represent the plant’s architecture. Modules in L-Strings can have parameters that give information about a plant organ. This includes the organ’s age, order in the branch or geometric attributes (e.g.  length, radius, angle, etc.). By applying geometric interpretation rules on an L-String using turtle geometry [1], a 3D scene of the plant can be generated. The virtual plant comprises of connected 3D shapes corresponding to the modules. In other words, an L-String can encode a plant’s geometry in addition to its architecture.
Our L-system models are implemented using the L-Py [2] language, that provides an extension of the Python programming language to design procedural models using the L-systems formalism. L-Py allows to create and manipulate L-system rules, handle the information present in an L-String, and generate 3D virtual plants from L-Strings.

### v2

@@ Abstract
Plants frequently contain numerous organs, organized in 3D branching systems defining the plant’s architecture. Reconstructing the architecture of plants from unstructured observations is challenging because of self-occlusion and spatial proximity between organs, which are often thin structures. To achieve the challenging task, we propose an approach that allows to infer a parameterized representation of the plant’s architecture from a given 3D scan of a plant. In addition to the plant’s branching structure, this representation contains parametric information for each plant organ, and can therefore be used directly in a variety of tasks. In this data-driven approach, we train a recursive neural network with virtual plants generated using a procedural model. After training, the network allows to infer a parametric tree-like representation based on an input 3D point cloud. Our method is applicable to any plant that can be represented as binary axial tree. We quantitatively evaluate our approach on Chenopodium Album plants on reconstruction, segmentation and skeletonization, which are important problems in plant phenotyping. In addition to carrying out several tasks at once, our method achieves results on-par with strong baselines for each task. We apply our method, trained exclusively on synthetic data, to 3D scans and show that it generalizes well.
@@ 1 Introduction
Plant phenotyping consists of quantifying how plant’s genotypes grow in their environment and has important applications in crop management. To automate this task, it is necessary to infer high-level information of plants from observations. This problem, which has been studied for decades [phenotyping2015], can be decomposed into a number of subtasks. High-level information includes the reconstruction of the 3D geometry of a plant, the segmentation of a plant into parts, and the extraction of a skeleton. Observations are typically 2D images or 3D scans. Automatically extracting this information has applications in plant phenotyping, e.g. [mirande2022graph, dobbs2023smart, ghahremani2021direct], and to use reconstructed 3D plants in content generation and virtual reality [TreePartNet21].
@@
Such an inference task is challenging as plants have complex structures due to their branching systems, which lead to strong self-occlusions and ambiguities. Existing works that analyze 2D or 3D plant observations therefore either focus on inverse modeling, where the goal is to find growth rules that allow to generate a given plant, or are task-specific. Inverse modeling is challenging and existing methods are limited to leafless branching structures [guo2020inverse, stava2014inverse]. Task-specific methods focus on 3D reconstruction e.g. [TreePartNet21], organ segmentation e.g. [mirande2022graph], or skeletonization e.g. [dobbs2023smart].
We propose a method to infer a parameterized representation given as input a 3D scan of a plant. This is achieved by learning a shape space of 3D plants that captures both the plant’s structure and the parametric 3D shape of all plant organs. The resulting representation can be directly used to reconstruct the 3D geometry of the plant with labels. This offers the key advantage of allowing to solve a variety of tasks, while allowing for complex plant architectures.
Learning shape spaces of plants is under-explored due to challenges arising from variation in both plant structure and organ shape. A recent work proposes a first solution [cheng2025demeter], while relying on large captured annotated datasets for training and requiring pre-segmented input for inference, both of which are costly and error-prone. In contrast, we propose a method trained purely on synthetic data and show its applicability to raw 3D scan acquisitions without annotations.
To learn a shape space of 3D plants, we leverage biologically inspired procedural models made of recursive rules that describe plant development. Procedural models inform both the design of our neural network that learns a shape space of 3D plants and the generation of large amounts of synthetic training data. We use Lindenmayer-systems (L-systems) to represent plants in a binary axial tree form [Prusinkiewicz1990TheAB]. In L-systems, each plant architecture is represented as a bracketed string of parameterized modules, the L-String. We use a recursive neural network to model the recursive nature of the L-string. Without loss of generality, we restrict plant representations to binary trees, which allows the use of recursive auto-encoders trained on (binary) L-Strings. In this way, we learn both branching structure and shape distribution of plant organs from synthetic data. To generalize to acquired 3D scans of plants, we learn a mapping from simulated 3D scan data to the learned shape space. For robustness, we simulate acquisition noise on the virtual plants.
We evaluate our method on Chenopodium album plants, an annual plant, in early growth stages. Chenopodium is an ideal plant for our tests as it is a small plant (as opposed to trees) while still displaying a complex branching architecture and has small leaves that partially occlude each other, but not excessively. In addition, a collection of real scans are readily available at different stages of development [mirande_2022_6962994]. We thus created for this plant a large dataset of 3D virtual plants represented as L-Strings, with associated 3D point clouds simulated with different types of acquisition noise.
We use our method to perform 3D reconstruction, skeleton extraction and plant organ segmentation. Our experiments show results on-par with strong baselines, and robust to noisy and partial inputs.
@@
  • An approach to infer a parameterized representation of a plant from a 3D scan based on a shape space of 3D plants learned using a recursive neural network.
  • 3D plant reconstruction, skeletonization and segmentation, with performance on-par with strong baselines.
  • A dataset of virtual plants, in the form of L-Strings, that represents instances of Chenopodium Album plants in early growth stages.
  • A demonstration that learning based exclusively on virtual plants can be generalized to fit real plants.
@@ 2 Related Work
Existing works to infer high-level information of plants belong to two main categories: inverse plant modeling and task-specific methods. Our work is between these categories, since it infers a parameterized plant representation, which includes the branching structure and 3D geometry of the plant. This infers a step of a procedural model and can be seen itself as a step towards inverse plant modeling.
A notable exception to our categorization is Demeter [cheng2025demeter], a work close in spirit to ours. Demeter learns a parametric plant model and demonstrates its use for reconstruction tasks. Unlike our method, Demeter requires manually annotated 3D scans for training and uses an elaborate three-step method for reconstruction, which requires pre-segmented input. In contrast, our method is trained on virtual data, which is cheaper to obtain, and directly infers a parametric representation without costly data pre-processing.
@@ 2.1 Inverse Plant Modeling
Procedural models generate plants following growth rules. In contrast, inverse plant modeling aims to find the rules that allow to generate a given plant. Guo et al[guo2020inverse] and Štáva et al[vst2010inverse] infer procedural modeling rules generated using an L-system from 2D images of leafless branching structures. Štáva et al[stava2014inverse] introduce a parametric procedural model in 3D specific to trees. None of these methods is applicable to annual plants with leaves. More recently, Lee et al[10.1145/3627101] used transformers to learn a generative model to represent L-system rules, without the aim of reconstructing a precise instance of a given real plant input. CropCraft [zhai2024cropcraftinverseproceduralmodeling] performs inverse procedural modeling taking as input a collection of 2D images, and not 3D scans.
@@ 2.2 Task Specific Methods
Work Recon- Skeletoni- Segment- Annual Direct
struction zation ation plants infer-
ence
Gonzalez et al[gonzalez2023tree]
Liu et al[TreePartNet21]
Du et al[du2019adtree]
Linvy et al[livny2010automatic]
Preuksakarn et al[plant-recon]
Parsad et al[prasad2022deep]
Dobbs et al[dobbs2023smart]
Chaudhury et al[Chaudhury2020.02.15.950519]
Yan et al[5246837]
Meyers et al[Meyer_2023]
Mirande et al[mirande2022graph]
Turgut et al[turgut2022segmentation]
Turgut et al[TURGUT2022138]
Wahabzada et al[2015article]
Li et al[LI2022243]
Li et al[psegnet2022]
Cheng et al[cheng2025demeter]
Ours
Table 1: Positioning of our method w.r.t. the ability to perform 3D reconstruction, skeletonization and segmentation, applicability on annual plants, and ability to perform inference directly without requiring pre-segmentation of the input data.
Existing plant phenotyping methods from 3D point clouds mainly focus on three tasks: 3D reconstruction, extracting a skeleton of the branching structure, and segmentation.
@@
While some works reconstruct 3D plants from a single 2D image [liu2025boxplant], we focus on reconstructing plant geometry from an input 3D point cloud. This 3D reconstruction problem is challenging as plants contain thin structures and self-occlusions.
In computer graphics, Gonzalez et al[gonzalez2023tree] reconstruct urban trees by computing a mesh representing the tree trunk, estimating the volume and density of the canopy, and filling the canopy with generated leaves. The branching structure of the plant is ignored. Another line of work focuses on branching structures of trees [du2019adtree, plant-recon, 5246837, TreePartNet21, 10.1145/1289603.1289610] by first extracting a skeleton and subsequently reconstructing a mesh per branch. Foliage is not reconstructed.
Closest to our work, Prasad et al[prasad2022deep] compare different implicit surface reconstruction algorithms on 3D point clouds of a single plant. We compare our method experimentally to the best performing method tested in this work.
@@
One line of work, discussed above, extracts skeletons as one part of a pipeline to reconstruct the plant [du2019adtree, plant-recon, 5246837, TreePartNet21, 10.1145/1289603.1289610]. Other approaches use optimization-based methods. Meyer et al[Meyer_2023] use a point contraction algorithm to extract skeletons of leafless trees. Graph-based approaches have shown to extract skeletal structures even in the presence of noise [10.1145/1289603.1289610, livny2010automatic]. A recent deep learning-based approach estimates the medial axis of a tree [dobbs2023smart]. Given a point cloud of a plant with an initial extracted skeleton, Chaudhury et al[Chaudhury2020.02.15.950519] refine the skeleton.
All discussed methods require leafless input. When leaves are present, skeletal structures are estimated inside the leaves, resulting in noisy output. Our method allows to extract skeletons with or without leaves. We experimentally compare our method to one successful approach from each category for which code is available [10.1145/1289603.1289610, livny2010automatic, Chaudhury2020.02.15.950519].
@@
Both learning and optimization-based techniques can segment a plant into organs. Mirande et al[mirande2022graph] propose a graph-based optimization approach with botanical knowledge refinement for semantic and instance segmentation. Wahabzada et al[2015article] present an unsupervised data-driven method. For supervised learning methods, a benchmark for plant organ segmentation [turgut2022segmentation] has been released based on the ROSE-X dataset [Dutagaci2020ROSEXAA]. More complex deep learning architectures designed for organ segmentation lead to accurate results [LI2022243, TURGUT2022138, psegnet2022]. We compare our method to the two generally applicable approaches [LI2022243, psegnet2022].
@@
Table 1 shows that only few works perform 3D reconstruction, skeletonization and segmentation [TreePartNet21, 5246837, cheng2025demeter]. Two of these [TreePartNet21, 5246837] are limited to big trees with trunks and foliage, and not applicable on small annual plants. The third method [cheng2025demeter] takes 2D images as input and performs segmentation as pre-processing, and is therefore not comparable to our method.
@@ 3 Background
This section provides background on plant modeling and recursive neural networks.
@@ 3.1 Plant Modeling
Plant modeling aims to find a mathematical model that describes the complex geometry and growth rules for a species of plants. Moreover, plant models respect biological rules to allow realistic simulation of the plant’s appearance and behavior in different environmental conditions [Godin.2005].
Procedural methods are among the most commonly used plant modeling approaches, including L-systems [Prusinkiewicz1990TheAB], the space colonization algorithm [space], and commercial tools such as SpeedTree [speedtree]. In our work, we use procedural plant models to generate training and test data.
Our implementation uses an L-system-based procedural approach [Prusinkiewicz1990TheAB] to generate plant data, a classical rule based plant modeling technique, that is particularly convenient due to the availability of tools such as L-Py [boudon2012py]. In L-systems, a plant is represented as a string of symbols, called L-String, possibly bearing geometrical or biological parameters. Similarly to natural or computer language grammars, an L-system consists of a set of rewriting rules that define how plant components represented by L-String symbols change as time proceeds, by specifying how symbols get replaced by combinations of other symbols [Prusinkiewicz1990TheAB, Godin.2005]. At each step, the rules replace the symbols in parallel, resulting in a new L-String representing the next plant state.
L-Strings of plants mainly consist of two types of symbols: modules and brackets. Modules can refer to various plant parts e.g. stems, leaves, flowers, etc. Opening and closing brackets indicate the start and the end of every branch in the plant. Two successive modules in an L-string have a parent-child relationship. Open brackets allow for a module to have more than one child, and for different children to have siblings relationships. However, each parent has at most one special child that corresponds to its successor on the same plant axis. Such L-Strings encode axial trees [Prusinkiewicz1990TheAB] that represent the plant’s architecture. Modules in L-Strings can have parameters that give information about a plant organ, which allows an L-String to encode a plant’s geometry in addition to its topology, defined as its structure.